OpenStudy (anonymous):

If an associate’s average sale to 5 customers is $125, and her sales to four of the customers are $200, $150, $125, and $75, what is the associate’s sale to the fifth customer? $75 $100 $50 $125 Not Here

sam (.sam.):

$75

OpenStudy (anonymous):

x/5= 125 where x = $200+ $150+ $125+ $75+ N where N is the 5 customer x=$550 + N ($550 + N)/5= $125 $550 + N= $ 625 N = $625- $550 = $75